OpenStudy (anonymous):

Can anyone help me answer this!?!?!?....Also its finding molarity not moles. Caluclate how many grams of solid Mg(NO3)2 and Sr(NO3)2 to dissolve in water to make your solutions. Specifically,write out the calculations to prepare 100.0ml of 0.1000M Mg(NO3)2 and 100.0 ml of 0.1000M Sr(NO3)2. --The Mg(NO3)2 supplied comes in the form of Mg(NO3)2 6H2O, use the formula weight 256.4 grams per mole in the calculation for Mg(NO3)2.

OpenStudy (anonymous):

u der??

OpenStudy (anonymous):

yes i am here :)

OpenStudy (anonymous):

i can help u too calculate as the que solution is lenghty i ll tell u what to do first acc. to dat u will do is it ok den we will cross chck our ans??

OpenStudy (anonymous):

okay that sounds wonderful! Thank you so much!

OpenStudy (anonymous):

have u written ur equation of reaction by balancing both sides

OpenStudy (anonymous):

no i am not sure how to do that...This pretty new to me...i am sorry

OpenStudy (anonymous):

its ok jusst tell me we have to find weight of both ubtance and molarity of equation rit

OpenStudy (anonymous):

ok firt we will calculate mass of mg(mg No3)2 formula mass/mw=v/22.4

OpenStudy (anonymous):

as mw=256.4 mass?? volume=1oo/1000 litre

OpenStudy (anonymous):

kelly r u der??

OpenStudy (anonymous):

yes i am writing this as well...i am following

OpenStudy (anonymous):

ok calculate it it must be approx=25.64 gram

OpenStudy (anonymous):

yes i got that too

OpenStudy (anonymous):

nw apply same formula for nxt compund u will get its mass

OpenStudy (anonymous):

as Sr(NO3)2=mw=160

OpenStudy (anonymous):

okay hold on please

OpenStudy (anonymous):

i came up with 241.6..i am not sure what i did wrong

OpenStudy (anonymous):

sorry listen in first compound i forgot to divide it by 22.4 so the ans of mg(NO3)2=1.15gm and the ans of Sr(NO3)2 is 0.75 gm

OpenStudy (anonymous):

for Sr(NO3)2 mw=160 v=100/1000 so , m/mw=v/22.4 m=vxmw/22.4 m=100x 160/22.4x 1000 m=0.75 gram

OpenStudy (anonymous):

you got it??

OpenStudy (anonymous):

okay yes i got it, thanks..and this to find the mass of the compound right?

OpenStudy (anonymous):

yess luk n=m/mw ................(1) n=v (in litre)/22.4............(2) from (1) and (2) we get m/mw=v/22.4

OpenStudy (anonymous):

nw for calculating molarity we have to write equation Mg(NO3)2 + Sr(NO3)2-------> Mg(NO3)2 + 6H20

OpenStudy (anonymous):

Mg(NO3)2 + Sr(NO3)2-------> Mg(NO3)2 + 6H20 +Sr

OpenStudy (anonymous):

jusst a doubt Mg(NO3)2, Sr(NO3)2 they are solid??

OpenStudy (anonymous):

yes they are solid

OpenStudy (anonymous):

which equation is the one you are using?

OpenStudy (anonymous):

the second one and for molarity we use mole in normality rit??

OpenStudy (anonymous):

moles= molarity x valency factor =normality = o.1000x 2

OpenStudy (anonymous):

okay, where did you get the 2?

OpenStudy (anonymous):

moarity=normality/volume as by above calculation u ll get normality and volume is 100/1000

OpenStudy (anonymous):

luk for valency factor calculate the charge present on Mg in Mg(NO3)2 by cross multiplication we get dis

OpenStudy (anonymous):

Okay I see, that makes sense

OpenStudy (anonymous):

now for molarity jut use d abv formula u ll get approx= 2

OpenStudy (anonymous):

as u also know MG i found in 2nd group so it has +2 charge or you heard o many times dat valency depend on the group num.

OpenStudy (anonymous):

your ques is finish any more doubt ??

OpenStudy (anonymous):

i am sorry heena, is there a way you can write this out like you did with the attachment you sent me...i am getting confused.

OpenStudy (anonymous):

ok in cross multiplication the chare get opposite like on NO3 it has -1 charge which goes to Mg and Mg has +2 charge which goes to NO3 o in compound we will write it like Mg(NO3)2

OpenStudy (anonymous):

yes i understand that...i am just lost when you are talking about morality.

OpenStudy (anonymous):

lolz.. its molarity ok no doubts now so solve it nd get ur ans bye..